OpenStudy (anonymous):

I need help on 7 math questions! I will fan+give 10 medals!!

OpenStudy (ribhu):

yes

OpenStudy (anonymous):

http://prntscr.com/67lam6

OpenStudy (anonymous):

what are the questions

OpenStudy (anonymous):

thats the first one ^

OpenStudy (anonymous):

I will give both of you the same reward if you guys help me :D

OpenStudy (ribhu):

24/20 = 1.2 i.e A

OpenStudy (anonymous):

my computer is being stupid so i cant see the problems

OpenStudy (ribhu):

@CREEPYMLPMASTER next one.

OpenStudy (anonymous):

alright let me get the next one and its okay red ill still reward you "D

OpenStudy (anonymous):

http://prntscr.com/67lbt6

OpenStudy (ribhu):

1323/27 = 49 i.e C

OpenStudy (anonymous):

so c?

OpenStudy (ribhu):

yaa

OpenStudy (anonymous):

http://prntscr.com/67lcru

OpenStudy (ribhu):

22/10 = 2.2 i.e A

OpenStudy (anonymous):

ok almost done :D http://prntscr.com/67ld9x

OpenStudy (ribhu):

next one please

OpenStudy (anonymous):

http://prntscr.com/67ld9x

OpenStudy (ribhu):

1.07

OpenStudy (ribhu):

next one.

OpenStudy (anonymous):

I am so sorry everything froze @ribhu